a school purchased five new basketballs for 149.95. at that rate, how much more money will it cost the school to have 12 new basketballs in all?
Answer by rfer(16322)

149.95/5=x/12
5x=149.95*12
5x=1799.40
x=1799.40/5
x=$359.88